OpenStudy (anonymous):

(1-i)^17/(1+isqur3)^9

OpenStudy (anonymous):

-1.81161

OpenStudy (anonymous):

its an ugly number

OpenStudy (helder_edwin):

write the complex numbers into its trigonometric form and use DeMoivre's law or use the binomial theorem on both

OpenStudy (anonymous):

it becomes squr2(cos 315+isin315)^17/2(cos60+isin60)

OpenStudy (anonymous):

^9

OpenStudy (anonymous):

2^7.5(cos(4815)+isin(4815))=0 ?

OpenStudy (anonymous):

does anything seem wrong?

OpenStudy (helder_edwin):

let's see

OpenStudy (helder_edwin):

\[ \large 1-i=\sqrt{2}(\cos(7\pi/4)+i\sin(7\pi/4)) \] \[ \large 1+i\sqrt{3}=2(\cos(\pi/3)+i\sin(\pi/3)) \] right?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

2=2^1 squr2=2^0.5

OpenStudy (anonymous):

17-9=8

OpenStudy (anonymous):

8-0.5=7.5

OpenStudy (anonymous):

thats how I thought

OpenStudy (helder_edwin):

then \[ \large (1-i)^{17}=\sqrt{2}^{17}[\cos(119\pi/4)+i\sin(119\pi/4)] \] \[ \large =2^8\cdot\sqrt{2}[\cos(14\cdot2\pi+7\pi/4)+i\sin(14\cdot2\pi+7\pi/4)] \] \[ \large =2^8\cdot\sqrt{2}[\cos(7\pi/4)+i\sin(7\pi/4)] \]

OpenStudy (anonymous):

I dont really understand...

OpenStudy (helder_edwin):

\[ \large (1+i\sqrt{3})^9=2^9[\cos(9\pi/3)+i\sin(9\pi/3)] \] \[ \large =2^9[\cos(3\pi)+i\sin(3\pi)]=2^9[\cos\pi+i\sin\pi] \]

OpenStudy (helder_edwin):

what don't u understand?

OpenStudy (anonymous):

the 119pi part

OpenStudy (helder_edwin):

do u know DeMoivre's theorem \[ \large (\rho(\cos\phi+i\sin\phi))^n=\rho^n(\cos(n\phi)+i\sin(n\phi)) \] ??

OpenStudy (anonymous):

Yes but how does that make it 119pi. I got 5355degrees and its not the same

OpenStudy (helder_edwin):

for \(z=1-i\) we have \(\rho=\sqrt{2}\) and \(\phi=7\pi/4\) right?

OpenStudy (anonymous):

So its wrong becouse I wrote in deggres and not pi? becouse 7pi/4=315 wich is what i got

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

Why do I need to write in pi? Thanks so much for helping me!

OpenStudy (helder_edwin):

it is usually better (actually customary) to use radians instead of degress when dealing with complex numbers

OpenStudy (helder_edwin):

u r welcome